A coin is placed 14.0 cm from the axis of a rotating turntable of variable speed. When the speed of the turntable is slowly increased, the coin remains fixed on the turntable until a rate of 31.0 rpm (revolutions per minute) is reached, at which point the coin slides off. What is the coefficient of static friction between the coin and the turntable?

Express answer using three different figures....
Answer #1

Solution: From the question we have:

A coin is placed 14.0 cm from the axis of a rotating turntable of variable speed

When the speed of the turntable is slowly increased,

the coin remains fixed on the turntable until a rate of 31.0 rpm (revolutions per minute) is reached

To find : What is the coefficient of static friction between the coin and the turntable?

=>The static coefficient of friction is equal to the centrifugal acceleration of the coin going around in a 14 cm radius circle at 31 rpm, divided by the acceleration due to gravity.
a = v^2/r
v = 14 cm x 2 x pi x 31 / min x 60 sec /min =45.42


a = v^2 / r= 45.42^2/14 = 147.39 cm / sec / sec

g = 9.8 x 100 cm / sec / sec
Coefficient = 147.39 / 980 = 0.150 ( note all the units cancel out) (two sig figs)